xcoledd1:

Help????!!! Algebra 2

xcoledd1:

xcoledd1:

Aratox:

Yeah... so here's the thing about Algebra... I passed both 1&2 but I don't think I'll be able to do it again

xcoledd1:

@aratox wrote:
Yeah... so here's the thing about Algebra... I passed both 1&2 but I don't think I'll be able to do it again
why

Aratox:

I didn't memorize it 💀... plus I guarantee you could find a calculator for this online :/

xcoledd1:

okay, calculator is blocked for me

xcoledd1:

@boldbrandon

xcoledd1:

@xcoledd1 wrote:
@boldbrandon
help

xcoledd1:

it's hard :skull:

xcoledd1:

can you help

xcoledd1:

...

xcoledd1:

It wants me to round

BoldBrandon:

4i (6 - 4i) = 4i 6 - 4i 4i = 24i - 16i^2 24i - 16(-1) = 24i + 16 16 + 24i (6 + 4i)(6 - 4i) = 6^2 - (4i)^2 = 36 - 16i^2 36 - 16(-1) = 36 + 16 = 52 52 16/52 + 24i/52 = 413 + 6i/13 4/13 + 6i/13: this should be your answer hehehe which idk i got it right but i am betting i sure did :D

xcoledd1:

@boldbrandon wrote:
4i (6 - 4i) = 4i 6 - 4i 4i = 24i - 16i^2 24i - 16(-1) = 24i + 16 16 + 24i (6 + 4i)(6 - 4i) = 6^2 - (4i)^2 = 36 - 16i^2 36 - 16(-1) = 36 + 16 = 52 52 16/52 + 24i/52 = 413 + 6i/13 4/13 + 6i/13: this should be your answer hehehe which idk i got it right but i am betting i sure did :D
thanks :3

BoldBrandon:

@xcoledd1 wrote:
@boldbrandon wrote:
4i (6 - 4i) = 4i 6 - 4i 4i = 24i - 16i^2 24i - 16(-1) = 24i + 16 16 + 24i (6 + 4i)(6 - 4i) = 6^2 - (4i)^2 = 36 - 16i^2 36 - 16(-1) = 36 + 16 = 52 52 16/52 + 24i/52 = 413 + 6i/13 4/13 + 6i/13: this should be your answer hehehe which idk i got it right but i am betting i sure did :D
thanks :3
yw :3

xcoledd1:

correct

BoldBrandon:

@xcoledd1 wrote:
correct
hell yes i am smart

xcoledd1:

@boldbrandon wrote:
@xcoledd1 wrote:
correct
hell yes i am smart
lol

BoldBrandon:

hehehe